# Weather-Driven Agricultural Decision-Making Using Digital Twins Under Imperfect Conditions

### v1

@@ 4.2.3. Model Training
We use partial grid search (bergstra2012random, ) to find out the optimized training hyper-parameters (see Table 2). We train all models with a batch size of 64 and apply early stopping to prevent over-fitting. We also use dropout regularization consistently across all the models, ranging from 0.1 to 0.3, depending on the architecture and robust generalization. Additionally, we standardize the weather parameters (i.e., features in the learning models) using a standard scaler to normalize them. The parameter labels are converted to one-hot encoded vectors for compatibility with the categorical cross-entropy loss function. All the models are saved as .keras format.
@@ 5.2. Inference Overheads and Scalability
In the next set of experiments (Fig. 5), we study the scalability of Cerealia. Specifically, we want to test if Cerealia can be used as a “digital twin as-a-service” model and serve various ag-tech vendors simultaneously. While the timing and memory overheads can be minimized by using powerful server-grade machines, we want to evaluate the effectiveness of our digital twin design in a relatively low-configured, low-cost device (such as our development platform, Jetson Orin, which has moderate CPU and GPU power and costs about \$2000). To test scalability, we create multiple instances of Cerealia using Docker containers (turnbull2014docker, ). Each container instance is a representation of the Cerealia digital twin “service” provided to the end-users. We deploy three configurations of $\{5,15,30\}$ containers and instantiate them using Docker Compose (version 1.29.2). Each container image is identical and configured with 1 CPU core and 4 GB of memory.
@@ 6.1. Case Study 1: Weather Data Imputation
Often, weather data becomes unavailable due to faults or breaches in the sensors or the entire station (boomgard2022machine, ). In addition, often, the end users may not have enough regional or site-specific data to use for analytics. In such cases, it is useful to predict sensor readings that can be “imputed” for missing or inconsistent results.
Our first case study shows the use of Cerealia to predict future weather measurements. Depending on how the models are trained, Cerealia can also mimic certain inconsistency patterns in sensor measurements. To predict sensor data for imputation, we use a generative model (C-RNN-GAN (mogren2016c, )), as this model is better suited for capturing short-term and long-term dependencies in time-series data. The blue lines in Fig. 6 show the ground truth data from the orchard (Quincy network), and the red dotted lines represent our predicted values obtained from the generative model. As shown in Fig. 6 and based on the error metrics presented in Table 5, the prediction results are pretty accurate.
@@ 6.2. Case Study 2: Fruit Surface Temperature Prediction
Heat stress to maturing fruit is a key concern for tree fruit growers. Localized weather-based fruit surface temperature (an indicator of fruit stress) can help in planning better mitigation strategies and ultimately reduce crop losses. To reduce fruit stress, growers use overhead cyclic rotating sprinklers, foggers, netting, and protectant sprays (racsko2012sunburn, ). However, precise data inputs are needed for the actuation and effective use of some of these mitigation techniques. Faulty or incorrect weather estimations can affect the initiation of such protective measures. We now show how Cerealia could be useful to predict fruit surface temperature so that growers can make informed decisions.
We use the aforementioned nine learning models to predict the apple surface temperatures from the weather measurements obtained from the Quincy network. The weather parameters for surface temperature prediction include canopy air temperature (C), wind speed (m/s), dew point (C), and solar radiation (W/m2). Past research shows fruit surface temperature can be estimated from these weather attributes (li2014modeling, ). Thanks to the AgWeatherNet team, we also obtained the “ground truth” apple temperature measurements. We study for both perfect and imperfect weather feeds. For imperfect cases, approximately 20% of the weather sensor feeds are faulty/inconsistent samples, and like before, we use equal noise splits (i.e., 5% for each class).
@@ 7. Related Work
Automating the modern agriculture sector is a promising research area. Zhai et al. (zhai2020decision, ) present a review of agricultural decision support systems by incorporating IoT, artificial intelligence, and big data. The use of digital twins for the agriculture sector is also studied. Peladarinos et al. (peladarinos2023enhancing, ) explore the role of digital twin in smart farming and emphasize their ability to create virtual replicas of agricultural systems by integrating sensors and remote imagery. The authors show that digital twins can support decision-making in irrigation, fertilization, and pest control using dynamic visualizations. Verdouw et al. (verdouw2021digital, ) propose a framework to integrate general systems control with IoT-based architecture to manage agricultural farm operations dynamically, which can enhance productivity, sustainability, and decision-making processes. An IoT-enabled digital twin pipeline is proposed (alves2019digital, ) that study integrates environmental data collection, cloud-based data processing, and real-time visualization to enhance resource management and optimize farming processes. None of the existing research focuses on investigating data anomalies in agricultural weather networks.
Digital twin technology is used in many cyber-physical domains for anomaly detection (castellani2020real, ; gao2021anomaly, ; xu2023digital, ). Castellani et al. (castellani2020real, ) propose a weakly-supervised anomaly detection technique for industrial systems using digital twins. The digital twin is used to generate synthetic data that represents normal operation conditions. The authors then use a set of labeled anomalies from real data to enhance detection accuracy. In a similar direction, Gao et al. (gao2021anomaly, ) introduce an anomaly detection framework for digital twin-driven cyber-physical systems, especially those that are vulnerable to sensor faults and model discrepancies. By integrating simulation data (from the digital twin) with real-time sensor data (from the physical plant), the proposed approach enables the detection of anomalies due to both sensor issues and model inaccuracies. LATTICE (xu2023digital, ) is a digital twin-based anomaly detection framework that uses curriculum learning to improve anomaly detection in cyber-physical systems. However, the above-mentioned techniques do not consider agricultural use cases. To the best of our knowledge, Cerealia is one of the first explorations for resiliency analysis of weather-driven agricultural systems.

### v2

@@ 4.2.3. Model Training
We use partial grid search (Bergstra and Bengio, 2012) to find out the optimized training hyper-parameters (see Table 2). We train all models with a batch size of 64 and apply early stopping to prevent over-fitting. We also use dropout regularization consistently across all the models, ranging from 0.1 to 0.3, depending on the architecture and robust generalization. Additionally, we standardize the weather parameters (i.e., features in the learning models) using a standard scaler to normalize them. The parameter labels are converted to one-hot encoded vectors for compatibility with the categorical cross-entropy loss function. All the models are saved as .keras format.
@@ 5.2. Inference Overheads and Scalability
In the next set of experiments (Fig. 5), we study the scalability of Cerealia. Specifically, we want to test if Cerealia can be used as a “digital twin as-a-service” model and serve various ag-tech vendors simultaneously. While the timing and memory overheads can be minimized by using powerful server-grade machines, we want to evaluate the effectiveness of our digital twin design in a relatively low-configured, low-cost device (such as our development platform, Jetson Orin, which has moderate CPU and GPU power and costs about \$2000). To test scalability, we create multiple instances of Cerealia using Docker containers (Turnbull, 2014). Each container instance is a representation of the Cerealia digital twin “service” provided to the end-users. We deploy three configurations of $\{5,15,30\}$ containers and instantiate them using Docker Compose (version 1.29.2). Each container image is identical and configured with 1 CPU core and 4 GB of memory.
@@ 6.1. Case Study 1: Weather Data Imputation
Often, weather data becomes unavailable due to faults or breaches in the sensors or the entire station (Boomgard-Zagrodnik and Brown, 2022). In addition, often, the end users may not have enough regional or site-specific data to use for analytics. In such cases, it is useful to predict sensor readings that can be “imputed” for missing or inconsistent results.
Our first case study shows the use of Cerealia to predict future weather measurements. Depending on how the models are trained, Cerealia can also mimic certain inconsistency patterns in sensor measurements. To predict sensor data for imputation, we use a generative model (C-RNN-GAN (Mogren, 2016)), as this model is better suited for capturing short-term and long-term dependencies in time-series data. The blue lines in Fig. 6 show the ground truth data from the orchard (Quincy network), and the red dotted lines represent our predicted values obtained from the generative model. As shown in Fig. 6 and based on the error metrics presented in Table 5, the prediction results are pretty accurate.
@@ 6.2. Case Study 2: Fruit Surface Temperature Prediction
Heat stress to maturing fruit is a key concern for tree fruit growers. Localized weather-based fruit surface temperature (an indicator of fruit stress) can help in planning better mitigation strategies and ultimately reduce crop losses. To reduce fruit stress, growers use overhead cyclic rotating sprinklers, foggers, netting, and protectant sprays (Racsko and Schrader, 2012). However, precise data inputs are needed for the actuation and effective use of some of these mitigation techniques. Faulty or incorrect weather estimations can affect the initiation of such protective measures. We now show how Cerealia could be useful to predict fruit surface temperature so that growers can make informed decisions.
We use the aforementioned nine learning models to predict the apple surface temperatures from the weather measurements obtained from the Quincy network. The weather parameters for surface temperature prediction include canopy air temperature (C), wind speed (m/s), dew point (C), and solar radiation (W/m2). Past research shows fruit surface temperature can be estimated from these weather attributes (Li et al., 2014). Thanks to the AgWeatherNet team, we also obtained the “ground truth” apple temperature measurements. We study for both perfect and imperfect weather feeds. For imperfect cases, approximately 20% of the weather sensor feeds are faulty/inconsistent samples, and like before, we use equal noise splits (i.e., 5% for each class).
@@ 7. Related Work
Automating the modern agriculture sector is a promising research area. Zhai et al. (Zhai et al., 2020) present a review of agricultural decision support systems by incorporating IoT, artificial intelligence, and big data. The use of digital twins for the agriculture sector is also studied. Peladarinos et al. (Peladarinos et al., 2023) explore the role of digital twin in smart farming and emphasize their ability to create virtual replicas of agricultural systems by integrating sensors and remote imagery. The authors show that digital twins can support decision-making in irrigation, fertilization, and pest control using dynamic visualizations. Verdouw et al. (Verdouw et al., 2021) propose a framework to integrate general systems control with IoT-based architecture to manage agricultural farm operations dynamically, which can enhance productivity, sustainability, and decision-making processes. An IoT-enabled digital twin pipeline is proposed (Alves et al., 2019) that study integrates environmental data collection, cloud-based data processing, and real-time visualization to enhance resource management and optimize farming processes. None of the existing research focuses on investigating data anomalies in agricultural weather networks.
Digital twin technology is used in many cyber-physical domains for anomaly detection (Castellani et al., 2020; Gao et al., 2021; Xu et al., 2023). Castellani et al. (Castellani et al., 2020) propose a weakly-supervised anomaly detection technique for industrial systems using digital twins. The digital twin is used to generate synthetic data that represents normal operation conditions. The authors then use a set of labeled anomalies from real data to enhance detection accuracy. In a similar direction, Gao et al. (Gao et al., 2021) introduce an anomaly detection framework for digital twin-driven cyber-physical systems, especially those that are vulnerable to sensor faults and model discrepancies. By integrating simulation data (from the digital twin) with real-time sensor data (from the physical plant), the proposed approach enables the detection of anomalies due to both sensor issues and model inaccuracies. LATTICE (Xu et al., 2023) is a digital twin-based anomaly detection framework that uses curriculum learning to improve anomaly detection in cyber-physical systems. However, the above-mentioned techniques do not consider agricultural use cases. To the best of our knowledge, Cerealia is one of the first explorations for resiliency analysis of weather-driven agricultural systems.
@@ 8. Conclusion
Acknowledgements.
We thank WSU AgWeatherNet team (Lav Khot and Srikanth Gorthi) for sharing weather station data. This research is partly supported by the U.S. National Science Foundation Award 2345653. Any findings, opinions, recommendations, or conclusions expressed in the paper are those of the authors and do not necessarily reflect the sponsor’s views.